OpenStudy (anonymous):

Can someone please help? I have to leave in like 30 minutes to pick my mom up from work and I still have 8 more questions. :) I wanna know this stuff for High School. :) @satellite73 @mathbrain99 @aric200 @iPwnBunnies @satellite73 @mathlover2014 @mathstudent55 @girlnotonfire @johnweldon1993 @Hero @OrangeMaster @RatzLover95 @Squirrels @Whitemonsterbunny17 @undeadknight26 @whpalmer4 @Yuba

OpenStudy (anonymous):

Explain? What is the value of n in the equation 3^2 × 3^3 = 3 n ? A. 3 B. 5 C. 6 D. 8

OpenStudy (aric200):

lego

OpenStudy (anonymous):

Please don't spam. I really need help. :(

hero (hero):

Hint: \(a^ba^c = a^{b + c}\)

OpenStudy (anonymous):

i think its 6

OpenStudy (anonymous):

So... B? 2+3=5....

OpenStudy (anonymous):

lolzies

OpenStudy (anonymous):

@Hero

hero (hero):

Correct @CrazyCountryGirl

OpenStudy (anonymous):

Thanks! That helped a lot. I'll have to right that down. Can you check my others? :) @Hero